OpenStudy (anonymous):

Suppose you choose a marble from a bag containing 2 red marbles, 5 white marbles, and 3 blue marbles. You return the first marble to the bag and then choose again. Find P(red and blue).

OpenStudy (anonymous):

P(red) = 2/10 = 1/5 P(blue) = 3/10 P(red and blue). = 1/5 * 3/10 = 3/50

OpenStudy (anonymous):

- this is the probability of picking a red first then a blue on second pick.

OpenStudy (anonymous):

if the order does not matter - that is you also allow blue first then red then the probability will be 2 * 3/50 or 3/25